A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1870–1879

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

30 November 2021 Version 1.00

Title word cross-reference

&c [Phi79].

1870 [Ano73a].

20th [Ano73a].

Abroad [Pag73]. Acclimatization [Nor75]. Account [Sto74, Mat79b]. Address [Ano75e]. American [Ano72, Ano73a, Ano74a, Ano75c, Ano76b, Ano77c, Ano78b, Ano79c, Edm72, Goo78b, Mat79b]. Annual [Ano75a, Ano77e, Ano78e, Ano79b]. Aquaculture [Wil75]. Aquaria [Mat79a]. Artificial [Mat73]. Association [Ano73a, Ano72, Ano74a, Ano75c, Ano76b, Ano77c, Ano78b, Ano78e, Ano79b, Ano79c, Mat79b].

B [Ano74c]. Bay [Mil78]. Blackford [Ano75g]. Breeding [Atk74, Sto74]. bronze [Phi79]. Brook [Cla79, Col72, Mat73]. Brook-Trout [Cla79]. Bucksport [Atk74].

California [Sto74]. Centennial [Ano77a]. Chesapeake [Mil78]. Chinese [Sal75]. collection [Phi79]. Commission [Bai74]. Committee [Ano74b, Ano75d, Ano76c, Ano77d, Ano78c, Ano79d]. Confinement [Mat78]. Constitution [Ano75b, Ano76a, Ano77b, Ano78a, Ano79a]. Cultural [Ano78b, Ano78e, Ano79b, Ano79c, Mat79b]. Culture [Bai73, Cli72, Mat73, Pag73, Sal75, Spr76, Sto72]. Culturist [Ano72, Ano73a, Ano74a, Ano75c, Ano76b, Ano77c, Gre74].

Day [Ano78d]. days [Mat79b]. December [Ano73a]. Depleted [Gre75].

E. [Ano75g]. early [Mat79b]. Eastern [Nor75, Sto74]. Eels [Roo78, Roo79]. Eggs [Dyk73, Sto78]. Eighth [Ano79b]. Enemies [McG79]. Establishments [Sto74]. Executive [Ano74b, Ano75d, Ano76c, Ano77d, Ano78c, Ano79d]. exhibiting [Phi79]. expenses [Mat79a]. Experiences [Gre74].

Features [Bla78]. Fecundation [Bel73]. Feeding [Mat78]. Fish [Ano72, Ano73a, Ano74a, Ano75c, Ano76b, Ano77c, Ano78b, Ano78e, Ano79b, Ano79c, Bai73, Bai74, Bel73, Bla78, Gre74, Gre76, Gre79, Hal74, Mat79b, Pag73, Phi79, Roc73, Sal75, Ste76, Sto74, Wil75, Wor75]. Fish-Market [Bla78]. Fish-Ways [Wor75]. Fisheries [Goo78b]. Fishery [Joh79]. Fishes [Goo78a, Mat78]. Fishways [Wor74]. Food [Cla79]. founders [Mat79b].

G [Ano75g]. General [Sto78]. Grayling [Nor75].

Habits [Roo79]. Hard [Ann79]. Hatching [Mil78]. Head [Mil78]. Hon. [Ano74c]. Hooks [Phi79].

Impregnation [Dyk73]. intentions [Mat79b]. Introduction [Edm72, Sto74].

Kiang [Sal75]. Kinds [Gre79].

Land [Bow72]. Land-Locked [Bow72]. Laws [Hal74]. Lobster [Joh79]. Locked [Bow72].

Management [Mat79a]. Market [Bla78]. McCloud [Sto74]. Meeting [Ano75a, Ano77a, Ano78e, Ano79b, Edm77, Ano77e]. Members [Ano73a, Ano74a, Ano75c, Ano76b, Ano77c, Ano78b, Ano79c]. Michigan [Nor75]. Migration [Goo78a].

National [Bai73]. Native [Roc73]. Natural [Mat73].

Obstructing [Mat75]. Officers [Ano73b, Ano74b, Ano75d, Ano76c, Ano77d, Ano78c, Ano79d]. Opening [Ano75e]. Operations [Sto74]. Organized [Ano73a].

Pacific [Sto74]. Packing [Sto78]. Peculiar [Bla78]. Pennsylvania [Wor74, Wor75]. plan [Mat79a]. Poisoning [Mat75]. Power [Roo78]. Practical [Gre74]. Practiced [Mat73]. Prehistoric [Phi79]. Preservation [Hal74]. President [Ano75e]. Proceedings [Ano78d, Ano72]. Propagation [Gre76, Ste76]. protect [Joh79]. Protection [Wil75]. Public [Mat79a].

Races [Col72]. Recollections [Mat79b]. reducing [Mat79a]. Remarks [Ano74c]. Report [Ano73c, Ano74d, Ano75f, Ano79e, Edm76, Sto78, Edm77]. Reproductive [Roo78, Roo79]. Resolutions [Ano75g]. River [Sto74]. Robert [Ano74c]. Roosevelt [Ano74c]. running [Mat79a].

Salmon [Atk74, Bow72, Edm72, Sto74, Sto78]. Salmon-Eggs [Sto78]. Samli [Sal75]. Second [Ano78d]. Seventh [Ano78e]. Shad [Cli72, Mil78, Sal75]. Shad-Hatching [Mil78]. Sixth [Ano77e]. Slope [Sto74]. Society [Edm77]. Spawning [Col72, Mat73]. Special [Edm77]. States [Sto74]. Statistics [Goo78b]. Stocking [Gre75, Gre79]. Subject [Sto78].

Teredo [Pag79]. their [Mat79a]. together [Sto74]. Transportation [Sto78]. Treasurer [Ano79e]. Trout [Ann79, Cla79, Col72, Dyk73, Mat73, McG79, Spr76, Sto72]. Tse [Sal75].

U.S. [Bai74]. United [Sto74]. Utah [Roc73].

Various [Gre79]. vs [Mat73].

Water [Ann79]. Waters [Edm72, Gre75, Gre79, Mat75, Mil78, Nor75, Sto74]. Ways [Wor75]. White [Ste76]. Whitebait [Bla79]. Work [Bai74, Mil78].

Yang [Sal75]. Yang-Tse-Kiang [Sal75].

References

Annin:1879:THW

[Ann79] James Annin, Jr. Trout in hard water. Transactions of the American Fisheries Society, 8(1):15–17, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1872:PAF

[Ano72] Anonymous. Proceedings of the American Fish Culturists’ Association. Transactions of the American Fisheries Society, 1(1):3–21, January 1872. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1873:MAF

[Ano73a] Anonymous. Members of the American Fish Culturists’ Association, Organized December 20th, 1870. Transactions of the American Fisheries Society, 2(1):33–34, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1873:O

[Ano73b] Anonymous. Officers. Transactions of the American Fisheries Society, 2(1):2, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1873:R

[Ano73c] Anonymous. Report. Transactions of the American Fisheries Society, 2(1):3–10, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1874:MAF

[Ano74a] Anonymous. Members of the American Fish Culturists’ Association. Transactions of the American Fisheries Society, 3(1):47–48, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1874:OEC

[Ano74b] Anonymous. Officers and Executive Committee. Transactions of the American Fisheries Society, 3(1):2, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1874:RHR

[Ano74c] Anonymous. Remarks of Hon. Robert B. Roosevelt. Transactions of the American Fisheries Society, 3(1):7–9, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1874:R

[Ano74d] Anonymous. Report. Transactions of the American Fisheries Society, 3(1):3–6, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1875:AM

[Ano75a] Anonymous. Annual meeting. Transactions of the American Fisheries Society, 4(1):6–11, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1875:C

[Ano75b] Anonymous. Constitution. Transactions of the American Fisheries Society, 4(1):5, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1875:MAF

[Ano75c] Anonymous. Members of the American Fish Culturists’ Association. Transactions of the American Fisheries Society, 4(1):42–43, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1875:OEC

[Ano75d] Anonymous. Officers, 1875–6 and Executive Committee. Transactions of the American Fisheries Society, 4(1):2, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1875:POA

[Ano75e] Anonymous. President’s opening address. Transactions of the American Fisheries Society, 4(1):12–14, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1875:R

[Ano75f] Anonymous. Report. Transactions of the American Fisheries Society, 4(1):3–4, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1875:RGB

[Ano75g] Anonymous. Resolutions of E. G. Blackford. Transactions of the American Fisheries Society, 4(1):11, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1876:C

[Ano76a] Anonymous. Constitution. Transactions of the American Fisheries Society, 5(1):17, January 1876. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1876:MAF

[Ano76b] Anonymous. Members of the American Fish Culturists’ Association. Transactions of the American Fisheries Society, 5(1):18–20, January 1876. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1876:OEC

[Ano76c] Anonymous. Officers, 1876–7 and Executive Committee. Transactions of the American Fisheries Society, 5(1):2, January 1876. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1877:CM

[Ano77a] Anonymous. Centennial meeting. Transactions of the American Fisheries Society, 6(1):10–45, January 1877. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1877:C

[Ano77b] Anonymous. Constitution. Transactions of the American Fisheries Society, 6(1):128, January 1877. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1877:MAF

[Ano77c] Anonymous. Members of the American Fish Culturists’ Association. Transactions of the American Fisheries Society, 6(1):129–131, January 1877. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1877:OEC

[Ano77d] Anonymous. Officers, 1877–8 and Executive Committee. Transactions of the American Fisheries Society, 6(1):2, January 1877. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1877:SAM

[Ano77e] Anonymous. Sixth Annual Meeting. Transactions of the American Fisheries Society, 6(1):46–127, January 1877. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1878:C

[Ano78a] Anonymous. Constitution. Transactions of the American Fisheries Society, 7(1):120, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1878:MAF

[Ano78b] Anonymous. Members of the American Fish Cultural Association. Transactions of the American Fisheries Society, 7(1):121–124, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1878:OEC

[Ano78c] Anonymous. Officers, 1878–9 and Executive Committee. Transactions of the American Fisheries Society, 7(1):2, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1878:SDP

[Ano78d] Anonymous. Second day’s proceedings. Transactions of the American Fisheries Society, 7(1):76, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1878:SAM

[Ano78e] Anonymous. Seventh Annual Meeting of the Fish Cultural Association. Transactions of the American Fisheries Society, 7(1):3–16, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1879:C

[Ano79a] Anonymous. Constitution. Transactions of the American Fisheries Society, 8(1):61, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1879:EAM

[Ano79b] Anonymous. Eighth Annual Meeting of the Fish Cultural Association. Transactions of the American Fisheries Society, 8(1):3–4, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1879:MAF

[Ano79c] Anonymous. Members of the American Fish Cultural Association. Transactions of the American Fisheries Society, 8(1):63–66, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1879:OEC

[Ano79d] Anonymous. Officers, 1879–80 and Executive Committee. Transactions of the American Fisheries Society, 8(1):2, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1879:TR

[Ano79e] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 8(1):62, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1874:SBB

[Atk74] Charles G. Atkins. Salmon breeding at Bucksport. Transactions of the American Fisheries Society, 3(1):24–30, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baird:1873:NFC

[Bai73] Spencer F. Baird. National fish culture. Transactions of the American Fisheries Society, 2(1):25–32, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baird:1874:WUF

[Bai74] Spencer F. Baird. Work of the U.S. Fish Commission. Transactions of the American Fisheries Society, 3(1):31–38, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1873:FF

[Bel73] Charles Bell. On the fecundation of fish. Transactions of the American Fisheries Society, 2(1):15–17, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackford:1878:PFF

[Bla78] Eugene G. Blackford. “Peculiar features of the fish-market”. Transactions of the American Fisheries Society, 7(1):77–87, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackford:1879:W

[Bla79] E. G. Blackford. Whitebait. Transactions of the American Fisheries Society, 8(1):11–15, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowles:1872:LLS

[Bow72] B. F. Bowles. Land-locked salmon. Transactions of the American Fisheries Society, 1(1):39–46, January 1872. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clapham:1879:FBT

[Cla79] Thomas Clapham. Food for brook-trout. Transactions of the American Fisheries Society, 8(1):4–6, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clift:1872:SC

[Cli72] Wm. Clift. Shad culture. Transactions of the American Fisheries Society, 1(1):21–28, January 1872. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1872:SRB

[Col72] A. S. Collins. Spawning races for brook trout. Transactions of the American Fisheries Society, 1(1):28–32, January 1872. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dykeman:1873:ITE

[Dyk73] J. R. Dykeman. On the impregnation of trout eggs. Transactions of the American Fisheries Society, 2(1):13–15, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edmonds:1872:ISA

[Edm72] M. C. Edmonds. The introduction of salmon into American waters. Transactions of the American Fisheries Society, 1(1):32–39, January 1872. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edmunds:1876:R

[Edm76] M. C. Edmunds. Report. Transactions of the American Fisheries Society, 5(1):3–7, January 1876. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edmunds:1877:RSM

[Edm77] M. C. Edmunds. Report: Special Meeting of the Society. Transactions of the American Fisheries Society, 6(1):3–9, January 1877. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1878:MF

[Goo78a] J. Browne Goode. Migration of fishes. Transactions of the American Fisheries Society, 7(1):27–67, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1878:SAF

[Goo78b] J. Browne Goode. Statistics of American fisheries. Transactions of the American Fisheries Society, 7(1):99–116, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1874:EPF

[Gre74] Seth Green. Experiences of a practical fish culturist. Transactions of the American Fisheries Society, 3(1):22–24, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1875:SDW

[Gre75] Seth Green. Stocking depleted waters. Transactions of the American Fisheries Society, 4(1):19–22, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1876:PF

[Gre76] Seth Green. Propagation of fish. Transactions of the American Fisheries Society, 5(1):8–13, January 1876. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1879:SWV

[Gre79] Seth Green. Stocking waters with various kinds of fish. Transactions of the American Fisheries Society, 8(1):22–27, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hallock:1874:LPF

[Hal74] Charles Hallock. Laws for the preservation of fish. Transactions of the American Fisheries Society, 3(1):45–46, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1879:LFH

[Joh79] S. M. Johnson. Lobster fishery and how to protect it. Transactions of the American Fisheries Society, 8(1):17–22, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1873:NVA

[Mat73] Fred Mather. Natural vs. artificial spawning as practiced in the culture of brook trout. Transactions of the American Fisheries Society, 2(1): 10–13, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1875:POW

[Mat75] Fred Mather. Poisoning and obstructing the waters. Transactions of the American Fisheries Society, 4(1):14–19, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1878:FFC

[Mat78] Frederick Mather. Feeding of fishes in confinement. Transactions of the American Fisheries Society, 7(1):67–75, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1879:MPA

[Mat79a] Frederick Mather. The management of public aquaria, with a plan for reducing their running expenses. Transactions of the American Fisheries Society, 8(1):46–51, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1879:RED

[Mat79b] Frederick Mather. Recollections of the early days of the American Fish Cultural Association, with an account of the intentions of its founders. Transactions of the American Fisheries Society, 8(1):55–60, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGovern:1879:NET

[McG79] H. D. McGovern. The new enemies of trout. Transactions of the American Fisheries Society, 8(1):6–11, January 1879. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Milner:1878:WSH

[Mil78] ?. Milner. The work of shad-hatching on the head waters of Chesapeake Bay. Transactions of the American Fisheries Society, 7(1):87–89, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norris:1875:AMG

[Nor75] Thaddeus Norris. On the acclimatization of the Michigan grayling in Eastern waters. Transactions of the American Fisheries Society, 4 (1):38–39, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:1873:FCA

[Pag73] George Shepard Page. Fish culture abroad. Transactions of the American Fisheries Society, 2(1):17–24, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:1879:T

[Pag79] George S. Page. Teredo. Transactions of the American Fisheries Society, 8(1):27–32, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1879:PFH

[Phi79] Barnet Phillips. Prehistoric fish hooks, exhibiting a collection of bronze hooks, &c. Transactions of the American Fisheries Society, 8(1):51–55, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rockwood:1873:NFU

[Roc73] A. P. Rockwood. The native fish of Utah. Transactions of the American Fisheries Society, 2(1):24–25, January 1873. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roosevelt:1878:RPE

[Roo78] Robert B. Roosevelt. Reproductive power of eels. Transactions of the American Fisheries Society, 7(1):90–99, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roosevelt:1879:RHE

[Roo79] Robert B. Roosevelt. The reproductive habits of eels. Transactions of the American Fisheries Society, 8(1):32–46, January 1879. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Salter:1875:FCY

[Sal75] G. H. Cotton Salter. Fish culture on the Yang-Tse-Kiang — the samli, or Chinese shad. Transactions of the American Fisheries Society, 4 (1):34–38, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sprout:1876:TC

[Spr76] A. B. Sprout. Trout culture. Transactions of the American Fisheries Society, 5(1):16, January 1876. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sterling:1876:PWF

[Ste76] E. Sterling. Propagation of white fish. Transactions of the American Fisheries Society, 5(1):13–15, January 1876. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1872:TC

[Sto72] Livingston Stone. Trout culture. Transactions of the American Fisheries Society, 1(1):46–56, January 1872. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1874:SBI

[Sto74] Livingston Stone. Salmon breeding: The introduction of Eastern fish into the waters of the Pacific Slope, together with an account of operations at the United States salmon breeding establishments on the McCloud River, California. Transactions of the American Fisheries Society, 3(1):9–22, January 1874. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1878:RGS

[Sto78] Livingstone Stone. Report on the general subject of the packing and transportation of salmon-eggs. Transactions of the American Fisheries Society, 7(1):16–27, January 1878. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilmot:1875:AFP

[Wil75] Samuel Wilmot. Aquaculture and fish protection. Transactions of the American Fisheries Society, 4(1):23–34, January 1875. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Worrall:1874:FP

[Wor74] James Worrall. The fishways of Pennsylvania. Transactions of the American Fisheries Society, 3(1):38–44, January 1874. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Worrall:1875:PFW

[Wor75] James Worrall. On the Pennsylvania fish-ways. Transactions of the American Fisheries Society, 4(1):40–41, January 1875. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).